OpenStudy (anonymous):

Find the surface area of a sphere with a radius of 60 m. Leave your answer in terms of pi @Directrix

Directrix (directrix):

The radius of the sphere is 60. A = 4 * π * r² A = 4 * π * (60)²

OpenStudy (anonymous):

14,400pi

Directrix (directrix):

I got 14,400π Good, you remembered the pi.

OpenStudy (anonymous):

yes